OpenStudy (anonymous):

Given the proportion (x+y)/3 = (y+z)/4 = (z+x)/5, determine the value of (xy+yz+zx)/x^2+y^2+z^2.

OpenStudy (anonymous):

\[\frac{ x+y }{ 3 }=\frac{ y+z }{ 4 }=\frac{ z+x }{ 5 }=k(say)\] x+y=3k (1) y+z=4k (2) z+x=5k (3) adding 2(x+y+z)=12 k x+y+z=6 k ....(4) subtract (1),(2),(3) from (4) z=3 k x=2 k y=k put the values of x,y,z and get the result.

OpenStudy (anonymous):

Thank you, it makes a lot more sense now!

OpenStudy (anonymous):

yw